OpenStudy (anonymous):

24/7 divide 3/2?? divide and simplify for me.. thanks

OpenStudy (anonymous):

48/21

OpenStudy (anonymous):

You'd cross multiply so...

OpenStudy (anonymous):

24/7 divided by 3/2 = 24/7 * 2/3 48/21 = 16/7

OpenStudy (anonymous):

You aren't supposed to solve the problems for them...

OpenStudy (yamyam70):